OpenStudy (anonymous):

Find the x-intercept of y = 4^x - 64.

OpenStudy (anonymous):

X intercept is when y=0 and x equal some value. 0=4^x-64 64=4^x ----I added 64 to both sides This is pretty straight forward from what I left off. Solve for x

OpenStudy (alexwee123):

make y=0 and solve out for x

OpenStudy (anonymous):

but look. a) (3,0) b) (64,0) c) (0.602,0) d) (1.806,0)

OpenStudy (zepp):

y = 4^x - 64 Set y as 0 and solve for x \[0 = 4^x - 64\]\[64 = 4^x\]\[\log_{4}64 = 3\]\[x=3\]

OpenStudy (anonymous):

Stop being lazy and solve for x stop trying to look at the answers first.

OpenStudy (anonymous):

... i just dont get it.

OpenStudy (anonymous):

You know how to solve 2x=1 for x?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

you did that by diving 2 to both sides and isolating x. You are doing the same thing in this case. 64= 4^x

OpenStudy (anonymous):

You want x by itself and you do this by ln both sides.

OpenStudy (anonymous):

but look at my answers. how is any of those answers get in that form?